# Hierarchy Pruning for Unseen Domain Discovery in Predictive Healthcare

### v2

@@ 4.3 More Quantitative Analysis
Figure 3: Effectiveness of Domain Discovery. The left figure shows the effect of the number of iteration on AUPRC in MIMIC-III dataset, and the right one shows comparison among variants upon UdonCare.
@@
Figure 4: Performance by Training Size. We evaluate drug recommendation on MIMIC-IV, and values on the x-axis indicate % of the entire training data. The dotted lines divide two ranges: [1, 10] and [10, 100].
Effect of Training Data Size. Next, inspired by Yang et al. (2023a); Jiang et al. (2024), we investigate how the volume of training data impacts model performance by conducting a comprehensive experiment in which the training set size ranges from 1% to 100%. Such a comparison is meaningful for examining how well models generalize with few domain samples. We evaluate drug recommendation on MIMIC-IV, since its complexity poses a challenging setting for prediction under varying EHR data sizes. All reported metrics are averaged over five independent runs. The results in Figure 4 indicate that all models show reduced performance in both AUPRC and F1-score when labeled data are scarce, particularly below 10% of the training set. However, UdonCare maintains a considerable edge over other baselines, suggesting that its co-training strategies effectively minimize information loss even when domain features are limited. Notably, MLDG lacks a certain level of resilience against data limitation, likely due to unique domain assignment, which might not work for the situation that existing patients with only few admissions (less than 3) on EHR datasets.
Table 4: Running Time Comparison of Drug Recommendation (seconds per epoch). Note that SLDG only use the most recent admissions for prediction.
@@
Runtime Analysis Lastly, we compare the training time of UdonCare with two other clinical DG baselines, ManyDG and SLDG, as shown in Table 4. All runtimes are measured on an NVIDIA L40S GPU. We find that an iterative training strategy effectively balances computational overhead and performance for the entire framework. We observe that UdonCare requires a training time comparable to SLDG, since both rely on iterative parameter updates that reduce the frequency of model adjustments during inference. ManyDG consumes more time than others, primarily because its domain assumption spawns numerous latent domains for subsequent computations.
5 Related Work
Domain Generalization (DG). DG is pursuing adjusted models which are specially designed to remove domain-covariate features from hidden representation (Muandet et al., 2013). A significant amount of work has been dedicated to solve performance drops on target domain across diverse scenarios like computer vision (Zhou et al., 2022; Ding et al., 2022), and they can be generally categorized as three different ways: (i) An intuitive way is to minimize the empirical source risk, either domain alignment (Ganin et al., 2016; Li et al., 2018b; Zhao et al., 2020) and invariant learning technique (Liu et al., 2021a; Zhang et al., 2022a; Wang et al., 2022) aim to convey little domain characteristics to acquire task-specific features. (ii) Contrastive learning (Kim et al., 2021; Jeon et al., 2021; Yao et al., 2022) becomes an alternative for data augmentation, studies employed the contrastive loss function to reduce the gap of representation distribution in one category. (iii) Meta-learning (Balaji et al., 2018; Dou et al., 2019) and ensemble-learning (Cha et al., 2021; Chu et al., 2022) approaches handle domain shifts through dynamic loss functions. However, they typically predefine either numerous or few domains in clinical settings (Wu et al., 2023), which motivates us to design a precise and efficient way to discover latent domains from learnable parameters.
DG in Clinical Prediction. Empirical evidence (Perone et al., 2019; Koh et al., 2021) has shown that EHR predictive models often suffer performance drops when transferred to new records with different data distributions. To address this, most existing works (Zhao et al., 2020; Zhang et al., 2021b) consider domain adaptation to handle potential domain shifts across multiple hospitals (Reps et al., 2022; Zhang et al., 2022b) and different time periods (Guo et al., 2022). Recently, a growing number of studies (Guo et al., 2022; Hai et al., 2024) consider model generalization by mitigating the patient-specific domain shifts, providing a more flexible alternative in more scenarios. For instance, Yang et al. (2023a) learns invariant features by treating each patient as a unique domain; Wu et al. (2023) develops a mixture-of-domain method to divide patients into latent domains by features of medical concepts. However, they primarily address domain categorization with simplifying heuristics such as linear dependencies (Li et al., 2020), while the potential of incorporating medical knowledge beyond EHR data remains unexplored in clinical DG problems.
@@ 6 Conclusion
This paper develops UdonCare, a novel framework for unseen domain discovery in predictive healthcare. Under the guidance of medical ontologies, our method discovers and iteratively adapts domain categorization. Extensive evaluations on two MIMIC datasets demonstrate that UdonCare outperforms state-of-the-art baselines across multiple tasks. For example, in mortality prediction, UdonCare surpasses other baselines by 4-8% in AUPRC; On readmission tasks, it also gains up to 6%, while drug recommendation has 5-10% improvements. Despite these gains, UdonCare remains comparable computational efficiency to other clinical DG baselines. These results demonstrate the good model generalization with knowledge-driven domain discovery in clinical practices. Future work can further reveal more benefits of medical knowledge in robust clinical predictions.
Ethics statement
We acknowledge that we have read and adhered to the conference Code of Ethics in the preparation and presentation of this work. In line with the principles of responsible stewardship, we are committed to upholding high standards of scientific excellence, honesty, and transparency in our research. We have conducted and presented this work with integrity, giving proper acknowledgment to the contributions of others and ensuring that our findings are reported accurately and reproducibly. We recognize the importance of minimizing potential harms and have reflected on the broader societal impacts of our research, including implications for human well-being and the natural environment. Consistent with the values of fairness and inclusivity, we support the equitable participation of all individuals in research and seek to promote accessibility and inclusiveness in both our methods and outcomes. We further respect the privacy and confidentiality of data that inform scientific discovery, and we endeavor to ensure that our work contributes positively to society, advances knowledge responsibly, and aligns with the long-term public good. At present, we do not identify any specific ethical concerns associated with this research.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. Details of the proposed method (Section 3.1 and Section 3.2) with training & inference procedures (Section 3.3) are provided, with pseudo code (Appendix B), implementation details (Appendix D) and complete results (Section 4.2 and Section 4.3) included. A complete description of the data processing steps is also provided in the supplementary materials. Furthermore, we supply anonymized source code of UdonCare in the supplementary materials to facilitate replication of our experiments.
Appendix
@@ Appendix A Related Work: Hierarchy-Aware Predictive Modeling
EHR prediction models can be broadly categorized into three types: RNN/CNN-based (Choi et al., 2017; Tong et al., 2025a), GNN-based (Lu et al., 2021; Jiang et al., 2024), and Transformer-based models (Shang et al., 2019a; Poulain & Beheshti, 2024; Tong et al., 2025b). Concretely, most graph-based studies (Choi et al., 2017; Shang et al., 2019a; Lu et al., 2021) utilize hierarchical medical classifications like ICD-9 (Organization et al., 1988) and ATC (Nahler & Nahler, 2009) to determine medical concept similarity by assuming diseases closer in the hierarchy share more characteristics, as reflected in similar embeddings. However, this method can be biased as it typically fails to capture complex relationships beyond simple parent-child links, such as complications or comorbidity, leading to sub-optimal predictions (Xu et al., 2023; 2024; Hu et al., 2024). Moreover, there is few research integrating these hierarchical structures with DG techniques, which could enhance model robustness across diverse healthcare settings. Properly leveraging hierarchical relationships in DG could improve the domain discovery process, ensuring models account for variance in disease manifestation across different patient demographics and regional practices. Thus, integrating hierarchy-aware modeling with DG approaches holds potential for developing more accurate and personalized predictive models in EHR, catering to the nuanced needs of global healthcare environments.
@@ Appendix B Pseudo Code for UdonCare
Since the training and inference phase has been explained in the main paper, we conduct our pseudo code by two consecutive phases:
@@
Input: EHR dataset $\mathcal{S}$ with patient’s data $\{\mathbf{x}^{(i)}\}_{i=1}^{N_{\mathrm{tr}}}$; Feature extractor $f_{\phi}$ with defined backbone (e.g. Transformer).
1:// When iteration $I=3$ and epochs $N=100$
2:for epoch $\in\{1,2,\dots,40\}$ do
3:  // Backbone Pathway
4:  Decode $\hat{y}_{p}\leftarrow d_{\eta}(f_{\phi,k}(x))$ with Equation 3.3;
5:end for
6:Obtain learned $\phi$; Initialize the hierarchy $\mathcal{H}$;
7:for iteration $\in\{1,2,3\}$ do
8:  // Hierarchy-Guided Domain Discovery
9:  Define & Update look-up table $\mathbf{M}$ via Step 1-4;
10:  Assign domain IDs $\mathcal{M}:\mathbf{x}\mapsto\mathbf{m}$ with Equation 2;
11:  // Self-Supervised Domain Encoding
12:  Initialize $g_{\theta}$; Pretrain $g_{\theta}$ by minimizing Equation 6;
13:  for epoch $\in\{1,2,\dots,20\}$ do
14:   // Domain Pathway
15:   Obtain patient embeddings $\mathbf{p}$ with Equation 1;
16:   Get domain features $\mathbf{r}$ by $g_{\theta}(\mathbf{m})$;
17:   Decompose invariant features $\mathbf{h}$ with Equation 7;
18:   Decode $\hat{y}_{p},\hat{y}_{h}$ with Equation 3.3;
19:   Minimize co-training loss $\mathcal{L}_{p},\mathcal{L}_{h}$ with Equation 3.3;
20:  end for
@@
Output: Trained models $f_{\phi}$, $g_{\theta}$, $d_{\eta}$, $q_{\xi}$; final prediction $\hat{y}_{h}$.
Appendix C Baselines
@@
Appendix D Implementation Details
Considering the common use of the Encoder-Decoder structure for clinical prediction, we adopt the Transformer (Vaswani et al., 2023) as the backbone feature extractor $f_{\phi}$ in UdonCare and all baselines. Specifically, we follow the implementation adapted from PyHealth (Yang et al., 2023b), consisting of three layers with a hidden size of 64, 4 attention heads, and a dropout rate of 0.2. The position encoding is applied across patient visits to capture temporal order. Diagnosis, treatment, and medication codes are embedded as separate feature keys using an embedding look-up table. The MLP classifiers used for both original and domain-invariant representations contain two hidden layers with sizes [64, 32], ReLU activations, and a dropout rate of 0.2, with task-specific output activations. We use the Adam optimizer for training, and all remaining hyperparameters follow the settings in PyHealth. For domain discovery, the score function uses $\alpha=0.5$, and the KL divergence loss coefficient $\lambda$ is set to 1.0 on MIMIC-III and 1.5 on MIMIC-IV. All models are trained for 100 epochs, and the best model is selected based on the AUPRC score monitored on the source validation set. We set the learning rate to $1\times 10^{-4}$ for $f_{\phi}$ and $5\times 10^{-5}$ for $g_{\theta}$, batch size to 32, iteration to 3, and self-supervised epoch to 30. We tune $\alpha$ and $\lambda$ based on validation performance. All experiments are conducted using Python 3.10 and PyTorch 2.3.1 with CUDA 12.4 on a server equipped with AMD EPYC 9254 24-Core Processors and NVIDIA L40S GPUs.
Appendix E Additional Experimental Results
E.1 Results with GAMENet Backbone
@@
E.2 Results with CGL Backbone
@@
Appendix F Domain Discovery with More Feature Keys
To examine whether latent domain discovery benefits from richer feature information, we further incorporate procedure and medication codes as additional keys in UdonCare. As shown in Table 7, the overall gains across prediction tasks are marginal and vary inconsistently across datasets. While certain metrics observe slight improvements, others remain unchanged or even decline. This suggests that introducing treatments and drugs into the domain partitioning process does not lead to stable enhancements. These findings are consistent with our conclusion that relying on the disease hierarchy (ICD-9-CM) provides a more effective and reliable basis for latent domain discovery, whereas incorporating additional feature keys yields only limited benefits.

### v3

@@ 5.4 More Quantitative Analyses
Figure 4: Effectiveness of Domain Discovery. The left figure shows the effect of the number of iteration on AUPRC, and the right one shows comparison among variants upon UdonCare.
@@
Table 3: Running Time Comparison of Drug Recommendation (seconds per epoch). Note that SLDG uses only the most recent admissions for prediction, following the original setting.
@@
Runtime Analysis Next, we compare the training time of UdonCare with two other clinical DG baselines, ManyDG and SLDG, as shown in Table 3. All runtimes are measured on an NVIDIA L40S GPU. We find that an iterative training strategy effectively balances computational overhead and performance for the entire framework. We observe that UdonCare requires a training time comparable to SLDG, since both rely on iterative parameter updates that reduce the frequency of model adjustments during inference. ManyDG consumes more time than others, primarily because its domain assumption spawns numerous latent domains for subsequent computations.
Figure 5: Performance by Training Size. We evaluate drug recommendation on MIMIC-IV, and the x-axis indicate % of the entire training data. The dotted lines divide [1, 10] and [10, 100].
Effect of Training Data Size. Lastly, inspired by (Yang et al., 2023a; Jiang et al., 2024), we investigate how the volume of training data impacts model performance by conducting a comprehensive experiment in which the training set size ranges from 1% to 100% to examine how well models generalize with few domain samples. We evaluate drug recommendation on MIMIC-IV, since its complexity poses a challenging setting for prediction under varying EHR data sizes. All reported metrics are averaged over five independent runs. The results in Figure 5 indicate that all models show reduced performance in both AUPRC and F1-score when labeled data are scarce, particularly below 10% of the training set. However, UdonCare maintains a considerable edge over other baselines, suggesting that its co-training strategies effectively minimize information loss even when domain features are limited. Notably, MLDG lacks a certain level of resilience against data limitation, likely due to unique domain assignment, which might not work for the situation that existing patients with only few admissions (less than 3) on EHR datasets.
@@ 6 Conclusion
In this work, we propose UdonCare, a novel framework for clinical domain generalization that discovers hierarchy-grounded domains with adaptive granularity. By pruning hierarchical concepts, our model addresses two unique challenges: the absence of domain IDs and the need for domain definitions that align with medical semantics. Extensive evaluations on multiple EHR datasets demonstrate that UdonCare outperforms existing baselines across diverse tasks, without incurring additional computational overhead. For example, in mortality prediction, UdonCare achieves a relative improvement of 11% in AUPRC over baselines. In the future, we plan to study multi-modal clinical prediction and more general ontology-rich problems.
Impact Statement
This paper presents work whose goal is to advance the field of domain adaptation in clinical prediction. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here. However, we extend the discussion of some potential directions beyond healthcare event prediction in Appendix L.
@@ Appendix A Related Work: Hierarchy-Aware Predictive Modeling
EHR prediction models can be broadly categorized into three types: RNN/CNN-based (Choi et al., 2017; Tong et al., 2025a), GNN-based (Lu et al., 2021; Jiang et al., 2024), and Transformer-based models (Shang et al., 2019a; Poulain and Beheshti, 2024; Tong et al., 2025b). Concretely, most graph-based studies (Choi et al., 2017; Shang et al., 2019a; Lu et al., 2021) utilize hierarchical medical classifications like ICD-9 (Organization and others, 1988) and ATC (Nahler and Nahler, 2009) to determine medical concept similarity by assuming diseases closer in the hierarchy share more characteristics, as reflected in similar embeddings. However, this method can be biased as it typically fails to capture complex relationships beyond simple parent-child links, such as complications or comorbidity, leading to sub-optimal predictions (Xu et al., 2023, 2024; Hu et al., 2024). Moreover, there is little research integrating these hierarchical structures with DG techniques, which could enhance model robustness across diverse healthcare settings. Properly leveraging hierarchical relationships in DG could improve the domain discovery process, ensuring models account for variance in disease manifestation across different patient demographics and regional practices. Thus, integrating hierarchy-aware modeling with DG approaches holds potential for developing more accurate and personalized predictive models in EHR, catering to the nuanced needs of global healthcare environments.
@@ Appendix B Pseudo Code for UdonCare
Since the training and inference phase has been explained in the main paper, we conduct our pseudo code by two consecutive phases, as shown in Algorithm 1:
@@
1:Input: EHR dataset $\mathcal{S}$ with patient’s data $\{\mathbf{x}^{(i)}\}_{i=1}^{N_{\mathrm{tr}}}$; Feature extractor $f_{\phi}$ with defined backbone (e.g. Transformer).
2:// When iteration $I=3$ and epochs $N=100$
3:for epoch $\in\{1,2,\dots,40\}$ do
4:  // Backbone Pathway
5:  Decode $\hat{y}_{p}\leftarrow d_{\eta}(f_{\phi,k}(x))$ with Equation (4.3);
6:end for
7: Obtain learned $\phi$; Initialize the hierarchy $\mathcal{H}$;
8:for iteration $\in\{1,2,3\}$ do
9:  // Hierarchy-Guided Domain Discovery
10:  Define & Update look-up table $\mathbf{M}$ via Step 1-4;
11:  Assign domain IDs $\mathcal{M}:\mathbf{x}\mapsto\mathbf{m}$ with Equation (2);
12:  // Self-Supervised Domain Encoding
13:  Initialize $g_{\theta}$; Pretrain $g_{\theta}$ by minimizing Equation (6);
14:  for epoch $\in\{1,2,\dots,20\}$ do
15:   // Domain Pathway
16:   Obtain patient embeddings $\mathbf{p}$ with Equation (1);
17:   Get domain features $\mathbf{r}$ by $g_{\theta}(\mathbf{m})$;
18:   Decompose invariant features $\mathbf{h}$ with Equation (7);
19:   Decode $\hat{y}_{p},\hat{y}_{h}$ with Equation (4.3);
20:   Minimize co-training loss $\mathcal{L}_{p},\mathcal{L}_{h}$ with Equation (4.3);
@@
22:end for
23:Output: Trained models $f_{\phi}$, $g_{\theta}$, $d_{\eta}$, $q_{\xi}$; final prediction $\hat{y}_{h}$.
Appendix C Notation Table
For clarity, we have attached a notation table here, summarizing all symbols used in the main paper, as shown in Table 4. The paper uses the generic notation $\mathbf{x}$ to denote patient input features. We use plain letters for scalars (e.g., $d,t,y$), boldface lowercase letters for vectors (e.g., $\mathbf{p},\mathbf{r},\mathbf{h}$), and boldface uppercase letters for matrices (e.g., $\mathbf{M}$). Euler script letters (e.g., $\mathcal{H},\mathcal{C}$) denote sets or hierarchies.
Table 4: Notations used in UdonCare.
Symbols Descriptions
$x^{(i)}$, $V_{t}^{(i)}$ EHR data of the $i$-th patient; the $t$-th visit of the $i$-th patient
$\mathbf{x}=\sum_{t=1}^{T}x_{t}$ Aggregated multi-visit representation used for domain discovery
$\mathcal{C}=\{c_{1},\ldots,c_{|\mathcal{C}|}\}$ Set of medical concepts (leaf-level ICD-9-CM codes)
$y^{(i)}\in\{0,1\}^{d}$ Ground-truth label vector for predictive tasks (binary or multi-label)
$f_{\phi,k}(\cdot)$, $\mathbf{p}_{k}$ Feature encoder for concept type $k$ (diagnosis, treatment, medication)
$\mathbf{p}_{k}=f_{\phi,k}(x_{k})$ The patient embedding for key $k$
$\mathbf{p}=\mathbf{p}_{1}\oplus\dots\oplus\mathbf{p}_{K}$ Concatenated patient-level embedding
$\mathcal{H}$ Concept hierarchy with $H$ levels (ICD-9-CM)
$n_{i}^{(h)}$ Node $i$ at hierarchy level $h$
$\mathrm{Desc}(n)$ Descendant leaf nodes of node $n$
$e_{i}$, $e_{i}^{(h)}\in\mathbb{R}^{h}$ Embedding of concept node or intermediate ontology node
$m:c_{i}\mapsto n_{i}^{(H)}$ Mapping from a medical concept to its leaf-level ontology node
$S(n)$ Score for pruning node $n$ based on purity, coverage, and depth
$\mathrm{cov}(n)$, $\mathrm{pur}(n)$, $\mathrm{dep}(n)$ Coverage, purity, and depth used in node scoring
$\alpha$, $\rho$ Hyperparameters controlling score balance and similarity threshold
$\mathbf{M}\in\{0,1\}^{N_{\mathrm{tr}}\times|C^{\prime}|}$ Domain-assignment matrix after pruning; encodes soft domain labels
$C^{\prime}$ Pruned code vocabulary representing discovered domains
$g_{\theta}(\cdot)$, $\mathbf{r}=g_{\theta}(\mathbf{m})$ Domain encoder producing latent domain representation $\mathbf{r}$
$\bar{\mathbf{p}}=\mathbb{E}[\mathbf{p}|\mathbf{m}]$ Mean patient embedding per domain, used for self-supervised alignment
$L_{r}$ Self-supervised loss (MSE + MMD) for domain encoder training
$\mathbf{h}=\mathbf{p}-\tilde{\mathbf{r}}$ Invariant feature obtained by removing domain-parallel components
$\tilde{\mathbf{r}}=\mathbf{r}\cdot\langle\mathbf{p}/\|\mathbf{r}\|,\,\mathbf{r}/\|\mathbf{r}\|\rangle$ Parallel projection of $\mathbf{p}$ onto domain direction $\mathbf{r}$
$d_{\eta}(\cdot)$, $\hat{y}_{p}$ Backbone prediction head producing $\hat{y}_{p}=d_{\eta}(\mathbf{p})$
$q_{\xi}(\cdot)$, $\hat{y}_{h}$ Invariant prediction head producing $\hat{y}_{h}=q_{\xi}(\mathbf{h})$
$\tilde{y}$ Averaged pseudo-label from $(\hat{y}_{p},\hat{y}_{h})$ for KL consistency
$L_{p}$, $L_{h}$ Mutual learning objectives combining CE loss and KL divergence
$\lambda$ KL divergence weight.
$\langle\cdot,\cdot\rangle$ Vector inner product
$\|\cdot\|$ Vector norm (Euclidean unless specified)
$W_{\text{cls}}(\cdot)$ the learned weight of a logistic regression task
Appendix D Baselines
@@
Appendix E Data Statistics
Table 5: Statistics of MIMIC-III and MIMIC-IV datasets.
MIMIC-III 6,497 42 2.66 13.06 4.54 33.71
MIMIC-IV 49,558 55 3.66 13.38 4.70 43.89
MIMIC-III and MIMIC-IV are widely used electronic health records in clinical prediction (Johnson et al., 2016, 2023). MIMIC-III covers ICU admissions from 2001 to 2012, while MIMIC-IV spans 2008 to 2019. To avoid overlapping time ranges with MIMIC-III, we only retain patients from the years 2013201320192019 in MIMIC-IV.
Appendix F Implementation Details
Considering the common use of the Encoder-Decoder structure for clinical prediction, we adopt the Transformer (Vaswani et al., 2023) as the backbone feature extractor $f_{\phi}$ in UdonCare and all baselines. Specifically, we follow the implementation adapted from PyHealth (Yang et al., 2023b), consisting of three layers with a hidden size of 64, 4 attention heads, and a dropout rate of 0.2. The position encoding is applied across patient visits to capture temporal order. Diagnosis, treatment, and medication codes are embedded as separate feature keys using an embedding look-up table. The MLP classifiers used for both original and domain-invariant representations contain two hidden layers with sizes [64, 32], ReLU activations, and a dropout rate of 0.2, with task-specific output activations. We use the Adam optimizer for training, and all remaining hyperparameters follow the settings in PyHealth. For domain discovery, the score function uses $\alpha=0.5$, and the KL divergence loss coefficient $\lambda$ is set to 1.0 on MIMIC-III and 1.5 on MIMIC-IV. All models are trained for 100 epochs, and the best model is selected based on the AUPRC score monitored on the source validation set. We set the learning rate to $1\times 10^{-4}$ for $f_{\phi}$ and $5\times 10^{-5}$ for $g_{\theta}$, batch size to 32, iteration to 3, and self-supervised epoch to 30. We tune $\alpha$ and $\lambda$ based on validation performance, which are 5e35e-3 and $0.5$, respectively. All experiments are conducted using Python 3.10 and PyTorch 2.3.1 with CUDA 12.4 on a server equipped with AMD EPYC 9254 24-Core Processors and NVIDIA L40S GPUs.
Appendix G Additional Experimental Results
G.1 Results with GAMENet Backbone
@@
G.2 Results with CGL Backbone
@@
G.3 Results of Cross-Institutional Distributional Shift
Following the setting used in SLDG (Wu et al., 2023), we extend our evaluation to the eICU dataset, where hospitals can be categorized into four groups based on their locations (Midwest, Northeast, West, and South). To quickly deploy our model on this dataset, we use the processing procedure of PyHealth. We evaluate the performance of our model under spatial domain shifts by treating the Midwest group as the target domain and the remaining groups as the source domain. We use diagnosis and treatment codes to construct patient records, as medications in eICU are not encoded with standard IDs. Table 8 reports the results for readmission and diagnosis prediction. We observe that UdonCare continues to demonstrate the superiority over all baselines across these two tasks.
Table 8: Performance under Cross-Institutional Distributional Shift on eICU.
Model Readmission Prediction Diagnosis Prediction
AUPRC AUROC w-F1 R@10
Oracle 22.74 (0.13) 69.29 (0.13) 63.18 (0.08) 79.05 (0.09)
Base 11.97 (0.08) 51.31 (0.06) 53.89 (0.05) 71.40 (0.05)
DANN 12.75 (0.16) 54.28 (0.14) 58.41 (0.18) 75.66 (0.24)
CondAdv 13.16 (0.18) 53.97 (0.12) 57.04 (0.23) 73.92 (0.17)
MLDG 14.82 (0.11) 54.25 (0.17) 58.13 (0.28) 75.31 (0.21)
IRM 15.73 (0.19) 58.31 (0.21) 59.22 (0.16) 76.46 (0.15)
PCL 15.14 (0.23) 58.06 (0.24) - -
ManyDG 17.53 (0.17) 60.20 (0.10) 60.61 (0.38) 76.94 (0.21)
SLDG 16.71 (0.21) 59.54 (0.18) - -
UdonCare 18.37 (0.18) 62.12 (0.11) 61.83 (0.14) 77.02 (0.11)
Appendix H Domain Discovery with More Feature Keys
To examine whether latent domain discovery benefits from richer feature information, we further incorporate procedure and medication codes as additional keys in UdonCare. As shown in Table 9, the overall gains across prediction tasks are marginal and vary inconsistently across datasets. While certain metrics observe slight improvements, others remain unchanged or even decline. This suggests that introducing conditions, treatments and drugs together into the domain partitioning process does not lead to stable enhancements. These findings are also consistent with our understanding, since we are more likely to get fine-grained domains: patients with the same disease may be split into different domains solely due to distinct medication regimens, which can weaken the benefits of ontology-guided grouping. Therefore, our conclusion is simply that using several ontologies at the same time is less efficient. Instead, we encourage practitioners to select the most appropriate ontology according to the task.
Note that, our reliance on the ICD-9-CM is not due to a limitation of our model, but because diseases are the most direct and required source of patient information. Almost all EHR prediction tasks rely on diagnostic codes as a primary input, and diseases exert strong influence on domain formation. Hence, the disease ontology serves as the most appropriate basis in the general clinical setting. It is also important to note that UdonCare is not restricted to disease ontologies. For instance, using a medication ontology (ATC classification) might be equally or even more appropriate in drug recommendation. Our model directly supports such substitution. In this paper, we want to align the used ontology across four tasks in our experiments.